Determine whether the graphs of the two lines are parallel.
y+3=9x
6x-y=-3
Are the graphs of the given equations parallel?
no or yes

Answer by Leaf W.(135)

NO.
To find whether two lines are parallel, put them into y = mx + b (slope-intercept) form, where the "m" is the slope. If the slopes are the same, the two lines are parallel.
y+3=9x
Subtract 3 from both sides: y=9x-3
y=9x-3 <==slope-intercept/y=mx+b form; "m" is 9, so the slope is 9
6x-y=-3
Subtract 6x from both sides: -y=-6x-3
Divide both sides by -1: y=6x+3
y=6x+3 <==slope-intercept/y=mx+b form; "m" is 6, so the slope is 6
The slopes are 9 and 6, which are not the same, so the lines are NOT PARALLEL.